# Efficient detection of chaos through the computation of the Generalized Alignment Index (GALI) by the multi-particle method

### v1

@@ 4.2 Hénon-Heiles model
We start by demonstrating that the computation of the GALI using the VM and MPM genuinely give the same results by considering initial conditions in Fig. 2. In Figs. 3 (a-b), we plot the time dependence of the two largest LEs [Eq. (4)] of the Lyapunov spectrum, $\lambda_{1}$ [blue curve] and $\lambda_{2}$ [green curve], for the initial condition of a regular orbit with $x_{1}=0$, $x_{2}=0.558$, $p_{2}=0$ and $p_{1}>0$ evaluated such that $\mathcal{H}=0.125$. Since we also require initial conditions for the tangent vectors, randomly selected coordinates drawn from a uniform distribution are used, and the deviation vectors are orthogonalized and rescaled to fit the desired magnitude for all tangent vectors $d_{0}=10^{-8}$. Note that we scale the deviation vector in the VM computations as good practice for comparison with the MPM.
@@
The LEs are computed using both the VM [Fig. 3(a)] and MPM [Fig. 3(b)]. In our simulations we integrate equations associated with each method up to time $t=10^{7}$ setting as integration time step $h=0.01$ which ensures that the relative energy error is bounded by $E_{r}\approx 10^{-11}$. Further, we fix the renormalization time $\tau=1$. For both methods, the computed LEs are decaying following the power law $\lambda_{1,2}\propto t^{-1}$ [black line in Fig. 3(a-b)]. This confirms that the studied orbit is regular. In addition, a divergence from the $t^{-1}$ power law of the computed LEs through of the MPM is building up from time $t\approx 10^{6}$ for which the magnitude of the $\lambda_{1,2}\approx 10^{-6}$. This divergence suggests a limit in the precision of the computation of the LEs using the MPM for regular orbits [29].
@@
Using the parameters above, we also compute the time evolution of the GALI. The numerical results of the time evolution of $\mbox{GALI}_{2}$ [blue curve], $\mbox{GALI}_{3}$ [green curve] and $\mbox{GALI}_{4}$ [red curve] are shown in Fig. 3(c) for the VM. In particular, we see that the VM is quite accurate in capturing the theoretical decay rates with slopes $-1$, $-2$ and $-4$ in the double logarithmic scale shown using the continuous, dashed, and dotted-dotted-dashed black lines in Figs. 3(c-d). The novelty in this work, is to perform the same computations of the GALIs, using the MPM and demonstrate the effectiveness of this approach. In Fig. 3(d), we show the time evolution of the $\mbox{GALI}_{2}$ [blue curve], $\mbox{GALI}_{3}$ [green curve] and $\mbox{GALI}_{4}$ [red curve] computed using the MPM with the same set of tangent vectors than in the VM counterpart. This panel displays a practically constant $\mbox{GALI}_{2}$ which remains comparable to the one obtained using the VM, blue curve of Fig. 3(c).
@@
On the other hand, for $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ an interesting trend is seen. At the early stage of the evolution, up to $t\lesssim 10^{3.5}$, the computed $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ are decaying following the same power laws as for the VM, black lines in Fig. 3(d). Then, for $t\gtrsim 10^{3.5}$ a levelling off to a practically constant value appears, taking the values of the computed GALI away from the expected power laws. As in case of the LEs computations, this moment marks the point from which the accumulations of numerical errors dominates the numerical values of the computed GALI.
Figure 4: 2D HH system. Same as in Fig. 3, but for a chaotic orbit with initial coordinates $x_{1}=0$, $p_{1}=0.42$, $x_{2}=-0.25$ and $p_{2}=0.$ on the Poincaré surface of section in Fig. 2. The filled black lines in (a-c) guide the eye with slope $-1$. On the other hand the filled, dashed, dotted-dotted-dashed lines in (c-d) guide the eye with slopes $-\lambda_{1}/\ln\left(10\right)$, $-2\lambda_{1}/\ln\left(10\right)$ and $-4\lambda_{1}/\ln\left(10\right)$ respectively with $\lambda_{1}\approx 0.0455$ at the final time of integration.
@@
For completeness we perform a similar analysis for a chaotic trajectory with initial coordinates on the PSS (Fig. 2) $x_{1}=0$, $x_{2}=-0.25$, $p_{2}=0$ and once again $p_{1}>0$ is selected such that $\mathcal{H}=0.125$. In Figs. 4(a-b) we plot the time evolution of the largest LEs when computed using both the VM and MPM respectively. As time evolves, the computed $\lambda_{1}$ for both methods show a striking resemblance, saturating to values $\lambda_{1}\approx 0.045$ at large time. On the other hand, the computed second LE $\lambda_{2}\propto t^{-1}$, black line in the same panels.
Figure 5: 2D HH system. Dependence of GALI on the norm of the initial deviation vectors $d_{0}$. Practical reliable regions of computation of the (a) GALI2, (b) GALI3 and (c) GALI4 of the HH using the VM (blue triangles) and MPM (red circles) methods. We averaged the GALIs over $1000$ sets of orthogonal tangent vectors of norm $d_{0}=\lVert\bm{w}_{0}\rVert=\lVert\bm{\delta}_{0}\rVert$ about the reference orbits with coordinates $x_{1}=0$, $x_{2}=-0.25$, $p_{1}=0.42$ and $p_{2}=0$. The final time of integration $T\approx 200$ time units.
@@
For this chaotic orbit, we also computed the temporal dependence of the associated GALI. In Figs. 4(c-d) we show the time evolution of $\mbox{GALI}_{2}$ [blue curve], $\mbox{GALI}_{3}$ [green curve] and $\mbox{GALI}_{4}$ [red curve]. The numerical results of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ are as predicted for the VM (Fig. 4(c)), with all the values of the computed GALIs following the theoretical decay rates $-0.02$, $-0.04$ and $-0.08$ respectively indicated by the continuous, dashed and dotted-dotted-dashed black lines in Figs. 4(c-d). We repeat the same simulations using the MPM and present the outcomes in Fig. 4(d). Overall, we see a good correspondence between the VM and MPM on numerical values of $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ for time $t\lesssim 200$. Nevertheless, for $t\gtrsim 200$, a divergence from the theoretical power laws can be seen with $\mbox{GALI}_{2}$ practically saturating at time $t=[500,~{}600]$.
We now turn to estimating the practical reliable regions of the computation of the GALI by exploring its dependence on the magnitude of separation vectors $d_{0}$. In Fig. 5, we plot $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ at $t\approx 200$ for the chaotic orbit used in Fig. 4. In addition, we calculate results averaged over $1000$ random sets of tangent vectors in order to mitigate the dependence of the computed GALI on the choice of the initial deviation vectors, (see e.g. Sec. 5.4.1.1 of Ref. [15]).
@@
The values of GALI computed using the VM are shown in blue triangles and the values obtained through the MPM in red circles. The numerically computed $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ using the VM are clearly independent of the magnitude of the initial deviation vectors in Figs. 5(a-c). For the MPM however, a non-trivial relation between the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ varying $d_{0}$ is found. Our results show $V-$like shapes with a global minimum value located at $d_{0}\approx 10^{-8}$ for all the presented GALI. This result is in agreement with the theoretical prediction [see e.g. Fig. 1(a)] which established the global minimum at $d_{0}\approx\varepsilon^{1/2}$ in a computer environment with precision $\varepsilon$. Consequently, in the case of double precision computation $\varepsilon\approx 10^{-16}$ the minimum occurs at $d_{0}\propto 10^{-8}$.
Figure 6: 2D HH system. Time evolution of the GALI: (a) $\mbox{GALI}_{2}$, (b) $\mbox{GALI}_{3}$ and (c) $\mbox{GALI}_{4}$ for an initial condition $x_{1}=0$, $x_{2}=0.558$, $p_{2}=-0.1$ and $p_{1}$ is taken such that $\mathcal{H}=0.125$, which corresponds to a regular orbit. For the MPM, we use three initial norms of the deviation vectors $d_{0}=10^{-5}$, $10^{-8}$ and $10^{-11}$ respectively the blue, green and red colored curve. The black curve in each panel represents the results obtained using the VM.
@@
To further probe the above observation on the numerical reliable regions above, we numerically evaluate the GALI using three different choices of the norm of the deviation vector $d_{0}$ for the regular orbit in Fig. 3. In Fig. 6, we show the time evolution of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ when computed using the MPM with initial norms of the separation vectors $d_{0}=10^{-5}$, $10^{-8}$ and $10^{-11}$ [blue, green and red curves respectively], and with the VM [black curves].
@@
In general the computation of the GALI of regular orbits is reliable in the limit of short-time for fairly small initial norms of the tangent vectors. Indeed, for all initial magnitudes of the separation vectors $d_{0}$ used in Fig. 6, the computation of $\mbox{GALI}_{2}$ in Fig. 6(a), $\mbox{GALI}_{3}$ in Fig. 6(b) and $\mbox{GALI}_{4}$ in Fig. 6(c) using the MPM practically overlap with the ones obtained using the VM up to $t\approx 10^{3}$. For larger values of the integration time $t\geq 10^{3}$, we expect the accumulation of numerical errors in case of the MPM to deviate the values of the computed GALIs from those obtained using the VM. Thus, based on the practical reliable regions of the MPM in Fig. 5, we expect the rate of accumulation of numerical errors to be the smallest for the case with $d_{0}=10^{-8}$. This is indeed what we observe in Fig. 6. The values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ computed with $d_{0}=10^{-8}$ [green curves in Fig. 6(a)-(c)] overlap with the results obtained with the VM for the longest integration times compared to the ones obtained using $d_{0}=10^{-5}$ and $d_{0}=10^{-11}$ [blue and red curves in Figs. 6(a)-(c)] .
@@
Let us now study the impact of the renormalization time during the computation of the GALI using the MPM. In order to reach our goal, we focus on the chaotic orbit used in Figs. 4 and 5. For our computations, we fix $d_{0}=10^{-9}$ and evolve the system’s orbit and its nearby orbits up to $t\approx 200$ with $h=0.01$. We record the final values of the $\mbox{GALI}_{2}$ in Fig. 7(a), $\mbox{GALI}_{3}$ in Fig. 7(b) and $\mbox{GALI}_{4}$ in Fig. 7(c) averaged over $1000$ sets of initially random orthogonal separation vectors, varying $\tau$ in the interval between $0.01$ to $100$. We find that the dependence of the GALI on the renormalization time $\tau$ can be split into two main regions. The first region corresponds to $\tau\lesssim 1$ for which the final values of the computed GALI remain practically constant at around $5\times 10^{-5}$, $2\times 10^{-10}$ and $1.2\times 10^{-18}$ for the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ respectively. On the other hand in the region where $\tau\gtrsim 1$, increasing the values of $\tau$ leads to a roughening and an overall increase in the final values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$. This results hints toward the appearance of large numerical errors in computing the GALI values.
@@
Figure 7: 2D HH system. Dependence of GALI2, GALI3 and GALI4 on (a-c) the renormalization time $\tau$ and (d-f) the integration time step $h$ for the chaotic orbit with $x_{1}=0$, $x_{2}=-0.25$, $p_{2}=0$ and $p_{1}$ is taken such that $p_{1}>0$ with $\mathcal{H}=0.125$. (a-c) We set the integration time step $h=0.01$. (d-f) We fixed the renormalization time $\tau=1$. In all computations, the final time of integration is $t=2\times 10^{2}$ with random initial deviation vectors of norm $d_{0}=10^{-9}$. The GALIs are averaged over $1000$ sets of random orthogonal deviation vectors.
@@
A similar behavior is also observed when computing the dependence of the computed GALI on the integration time step $h$. For these numerical simulations, we kept the same set up for the numerical integration as above, setting $\tau=1$ and varying the integration time step $h$. The obtained results are depicted in Figs. 7(d)-(f) for the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ respectively. We see that the dependence of the final value of the GALI on $h$ can also be split into two main regions, separated by $h\approx 10^{-1}$. For $h<0.1$, these GALI values are constant for all $h$, while in $h>0.1$ we obtain different final values of the GALIs for each time step $h$. The striking results here is that in term of the relative energy error, $h=0.1$ corresponds to $E_{r}\approx 10^{-8}$, Fig. 8. It follows that numerical simulations with relative energy error $E_{r}>10^{-8}$ are in general inaccurate. Such a precision could be easily achieved using high order numerical integration schemes. It is worth pointing out the slope of the $E_{r}$ against $h$ relation which gives $4$ in Fig. 8 matching the order of the integration method, see also Sec. 3.2.
@@
Figure 8: 2D HH system. Dependence of the upper bound of the time evolution of the relative energy error $E_{r}$ on the integration time step $h$. For each point in the graph the initial condition ($x_{1}=0$, $x_{2}=-0.25$, $p_{2}=0$ and $p_{1}>0$ ($\mathcal{H}=0.125$), is integrated up to $t=200$ with the $ABA864$ symplectic integration scheme of order $4$.
Figure 9: 2D HH system. Regions of different dynamical behavior on the PSS of the HH at $x_{1}=0$ with $\mathcal{H}=0.125$ [Eq. (38)] computed through (a) the MLE at $t=1000$, (b) $\mbox{GALI}_{2}$ at $t=500$. (c) Regions of different values of the time $t$ needed for $\mbox{GALI}_{3}$ to become less than $10^{-8}$ on the same PSS section. In all computations $h=0.01$, $\tau=1$ and $d_{0}=10^{-8}$. In all panels, $50$ sets of initial orthogonal separation vectors were used for each initial condition.
@@
What about exploring the global dynamics of a dynamical system using the GALI method with the MPM? In Fig. 9(a-b), we show the colormap of the PSS at $x_{1}=0$ of Fig. 2 made of $700\times 600$ grid points. The remaining coordinate of each initial conditions on the ($x_{2}$, $p_{2}$) phase plane of Fig. 9 is set such that $p_{1}>0$ with $\mathcal{H}=0.125$. Each generated orbit on the $(x_{2},p_{2})$ phase plane is colored according its MLE value at $t=1000$ [Fig. 9(a)] and $\mbox{GALI}_{2}$ value at $t=500$ [Fig. 9(b)]. In addition the results are averaged over $50$ configurations of sets of random orthogonal tangent vectors in order to mitigate their dependence on the choice of the initial conditions for the tangent vectors. Both chaos indicators display the same qualitative dynamical behavior in terms of characterizing the stability of orbits in the phase space.
More specifically, Figs. 9(a-b), return the same phase plane structure, that is to say the regular [yellow colored points] regions and chaotic seas [purple colored points] are practically the same in both panels and similar to what is revealed by the PSS of Fig. 2. In particular, we see that several small regular islands in the chaotic regions are well capture by the MPM, along with the so-called ‘sticky orbits’ at the interfaces of regular islands with chaotic seas with a mixed blend of colors [e.g., yellow purple]. Here the difference between the two numerical simulations resides on the computational time. Indeed, it took $t=35.7$ and $t=28.2$ hours respectively for the simulations of Fig. 9(a) and Fig. 9(b) to complete [65].
To the question whether we can reach the same accuracy analyzing the HH phase space with GALI of higher orders, the answer is yes. First we remind that the GALI of higher orders decrease for both regular and chaotic orbits. It follows that we need to rely on a different approach than recording the final value of the GALI as above. In order to characterize the global dynamics of the HH, we register the time it took the GALI to reach a particular threshold [15] since the GALI values tend to decay faster for chaotic orbits compared to regular ones. In Fig. 9(c), we show the time it took $\mbox{GALI}_{3}$ to reach the threshold $10^{-8}$. This time is averaged over $50$ different sets of orthogonal tangent vectors at each point once again to mitigate the dependence on the choice of the initial tangent vectors.

### v2

@@ 4.2 The Hénon-Heiles model
We start our study by demonstrating that the computation of the GALIs using the VM and the MPM genuinely give the same results by considering specific initial conditions in Fig. 2. In Figs. 3 (a-b), we plot the time dependence of the two largest LEs $\lambda_{1}$ (blue curves) and $\lambda_{2}$ (green curves), for the regular orbit with initial conditions $x_{1}=0$, $x_{2}=0.558$, $p_{2}=0$ and $p_{1}>0$ evaluated such that $\mathcal{H}=0.125$. Since we also require initial conditions for the tangent vectors, randomly selected coordinates drawn from a uniform distribution are used, and the created deviation vectors are orthogonalized and rescaled to fit the desired magnitude for all tangent vectors $d_{0}=10^{-8}$. Note that we scale the deviation vectors in the VM computations as good practice for comparison with the MPM.
Figure 3: 2D HH system. Time evolution of (a-b) the two largest finite-time LEs $\lambda_{1}$ (blue curves) and $\lambda_{2}$ (green curves), and (c-d) the $\mbox{GALI}_{2}$ (blue curves), $\mbox{GALI}_{3}$ (green curves) and $\mbox{GALI}_{4}$ (red curves) for a regular orbit with initial conditions $x_{1}=0$, $p_{1}>0$, $x_{2}=0.558$ and $p_{2}=0$ of the HH with $\mathcal{H}=0.125$. The results are obtained by (a), (c) the VM and (b), (d) the MPM. The continuous black lines in (a-b) guide the eye for slope $-1$, while the continuous, dashed, and dashed-dotted ones in (c-d) have slopes 00 (horizontal lines), $-2$ and $-4$ respectively. All the calculations are performed for $d_{0}=10^{-8}$.
@@
The finite-time LEs are computed using both the VM [Fig. 3(a)] and the MPM [Fig. 3(b)]. In our simulations we integrate equations associated with each method up to a final time $t=10^{7}$, setting the integration time step to $h=0.01$, which ensures that the relative energy error [Eq. (41)] is upper bounded by $E_{r}\approx 10^{-11}$. Furthermore, we fix the renormalization time to $\tau=1$. For both methods, the computed LEs are decaying following the power law $\lambda_{1,2}\propto t^{-1}$ [black lines in Figs. 3(a-b)]. This behavior confirms that the studied orbit is regular. In addition, a divergence from the $t^{-1}$ power law of the computed LEs through of the MPM is observed after $t\approx 10^{6}$ for which the magnitudes of the LEs are $\lambda_{1,2}\approx 10^{-6}$. This divergence suggests a limit in the precision of the computation of the LEs using the MPM for regular orbits [31].
@@
Using the same model parameters, we also compute the time evolution of the GALIs. The numerical results of the time evolution of $\mbox{GALI}_{2}$ (blue curve), $\mbox{GALI}_{3}$ (green curve) and $\mbox{GALI}_{4}$ (red curve) are shown in Fig. 3(c) for the VM. In particular, we see that the VM is quite accurate in capturing the theoretical decay rates with slopes $-2$ and $-4$, for the GALI3 and the GALI4 respectively, which are indicated by the dashed and dashed-dotted black lines in Fig. 3(c), as well as the practical constancy of GALI2 [horizontal black line in Fig. 3(c)]. The novelty of our work, is that we perform the same computations of the GALIs, using the MPM and demonstrate the effectiveness of this approach.
@@
In Fig. 3(d) we show the time evolution of the $\mbox{GALI}_{2}$ (blue curve), $\mbox{GALI}_{3}$ (green curve) and $\mbox{GALI}_{4}$ (red curve) computed using the MPM with the same set of tangent vectors used in the VM computations of Fig. 3(c). There we see a practically constant $\mbox{GALI}_{2}$ which remains similar to the one obtained using the VM [blue curve of Fig. 3(c)]. On the other hand, for both the $\mbox{GALI}_{3}$ and the $\mbox{GALI}_{4}$ an interesting behavior is seen. At the early stage of the evolution, up to $t\lesssim 10^{3.5}$, the computed $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ are decaying following the same power laws [respectively denoted by the dashed and dashed-dotted black lines in Fig. 3(d)] as for the VM. Then, for $t\gtrsim 10^{3.5}$, a leveling off to a practically constant value appears, taking the values of the computed GALIs away from their expected power laws. As in the case of the LEs computations, that time marks the point from which the accumulations of numerical errors dominates the numerical values of the computed GALIs.
@@
We also perform a similar analysis for a chaotic orbit of the HH system, with initial conditions on the $x_{1}=0$ PSS of Fig. 2 $x_{2}=-0.25$, $p_{2}=0$, while, once again, $p_{1}>0$ is selected such that $\mathcal{H}=0.125$. In Figs. 4(a-b) we plot the time evolution of the two largest finite-time LEs computed using both the VM and MPM respectively. As time evolves, the computed $\lambda_{1}$ (blue curves) for both methods show a striking resemblance, saturating to values $\lambda_{1}\approx 0.045$ at large time. On the other hand, the computed second LE (green curves) show a power law decay $\lambda_{2}\propto t^{-1}$, which is denoted by the black lines in both panels.
Figure 4: 2D HH system. Similar to Fig. 3, but for a chaotic orbit with initial coordinates $x_{1}=0$, $p_{1}=0.42$, $x_{2}=-0.25$ and $p_{2}=0$. The continuous black lines in (a-b) guide the eye for slope $-1$. The continuous, dashed and dashed-dotted lines in (c-d) indicate slopes $-\lambda_{1}/\ln\left(10\right)$, $-2\lambda_{1}/\ln\left(10\right)$ and $-4\lambda_{1}/\ln\left(10\right)$ respectively, with $\lambda_{1}=0.045$.
@@
For this chaotic orbit, we also computed the temporal dependence of the associated GALIs. In Figs. 4(c-d) we show the time evolution of the $\mbox{GALI}_{2}$ (blue curves), $\mbox{GALI}_{3}$ (green curves) and $\mbox{GALI}_{4}$ (red curves). The numerical results of all GALIs for the VM are in accordance to the theoretical expectations [Eq. (8)], as the values of $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ follow exponential decays with exponents $-\lambda_{1}/\ln\left(10\right)\approx-0.02$, $-2\lambda_{1}/\ln\left(10\right)\approx-0.04$ and $-4\lambda_{1}/\ln\left(10\right)\approx-0.08$ respectively, for $\lambda_{1}=0.045$, which are indicated by the continuous, dashed and dashed-dotted black lines in Fig. 4(c) [these exponential decays are also indicated in the same way in Fig. 4(d)]. The outcomes of the same computations using the MPM are presented in Fig. 4(d). Overall, we see a good agreement between the VM and the MPM for the numerical values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ for times $t\lesssim 200$. Nevertheless, for $t\gtrsim 200$ a divergence from the theoretical power laws can be seen with $\mbox{GALI}_{2}$ practically saturating at times $t\gtrsim 500$.
@@
We now estimate the practical reliable regions of the computation of the GALIs by exploring the indices’ dependence on the magnitude of separation vectors $d_{0}$. In Fig. 5, we plot the values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ at $t=200$ for the chaotic orbit considered in Fig. 4, as a function of $d_{0}$. We note that the presented results are obtained by averaging over $1000$ random sets of initial tangent vectors in order to mitigate the dependence of the computed GALIs on the choice of the initial deviation vectors, (see e.g. Sect. 5.4.1.1 of [18]). The values of the GALIs computed using the VM are indicated by blue triangles, while the values obtained through the MPM are shown by red circles. The numerically computed $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ values using the VM are clearly independent of the magnitude of the initial deviation vectors in Fig. 5. For the MPM however, a non-trivial relation between the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$, $\mbox{GALI}_{4}$ values and $d_{0}$ is seen. Our results show $V-$like shapes with a global minimum value located at $d_{0}\approx 10^{-8}$ for all presented GALIs. This result is in agreement with the theoretical prediction [see Fig. 1(a)] which established the global minimum at $d_{0}\approx\varepsilon^{1/2}$ in a computer environment with precision $\varepsilon$. Consequently, in the case of double precision computations for which $\varepsilon\approx 10^{-16}$ the minimum occurs at $d_{0}\approx 10^{-8}$.
Figure 5: 2D HH system. Dependence of the GALIs on the norm $d_{0}$ of the initial deviation vectors. Practical reliable regions of the computation of the (a) GALI2, (b) GALI3 and (c) GALI4 for the HH model using the VM (blue triangles) and the MPM (red circles). The presented GALI values have been averaged over $1000$ sets of initial orthogonal tangent vectors of norm $d_{0}=\lVert\bm{w}_{0}\rVert=\lVert\bm{\delta}_{0}\rVert$ about the reference chaotic orbit with coordinates $x_{1}=0$, $x_{2}=-0.25$, $p_{1}=0.42$ and $p_{2}=0$. The final time of integration is $t=200$ time units.
@@
To further probe these findings on the numerical reliable regions, we numerically evaluate the GALIs using three different choices of the norm of the deviation vector $d_{0}$ for the regular orbit of Fig. 3. In Fig. 6, we show the time evolution of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ when computed using the MPM with initial norms of the separation vectors $d_{0}=10^{-5}$, $10^{-8}$ and $10^{-11}$ (blue, green and red curves respectively), and with the VM (black curves).
Figure 6: 2D HH system. Time evolution of the (a) $\mbox{GALI}_{2}$, (b) $\mbox{GALI}_{3}$ and (c) $\mbox{GALI}_{4}$ for the regular orbit of Fig. 3. For the application of the MPM we use three different values for the initial norms of the deviation vectors: $d_{0}=10^{-5}$, $10^{-8}$ and $10^{-11}$ (respectively blue, green and red colored curves). The black curve in each panel represents results obtained using the VM.
@@
From the results of Fig. 6 we see that, in general, the computation of the GALIs for regular orbits is reliable for short-time evolution of fairly small initial norms of the tangent vectors. Indeed, for all initial magnitudes of the separation vectors $d_{0}$ used in Fig. 6, the computations of the $\mbox{GALI}_{2}$ [Fig. 6(a)], $\mbox{GALI}_{3}$ [Fig. 6(b)] and $\mbox{GALI}_{4}$ [Fig. 6(c)] using the MPM practically overlap with the ones obtained using the VM up to $t\approx 10^{3}$. For larger values of the integration time ($t\gtrsim 10^{3}$) we expect the accumulation of numerical errors in the case of the MPM to result in values of the computed GALIs which deviate from those obtained using the VM. Thus, based on the practical reliable regions of the MPM in Fig. 5, we expect the rate of accumulation of numerical errors to be the smallest for the case with $d_{0}=10^{-8}$. This is indeed what we observe in Fig. 6. The values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ computed with $d_{0}=10^{-8}$ [green curves in Fig. 6] overlap with the results obtained with the VM for longer integration times compared to the ones obtained using $d_{0}=10^{-5}$ and $d_{0}=10^{-11}$ [respectively blue and red curves in Fig. 6].
@@
Let us now investigate the impact of the renormalization time $\tau$, on the computation of the GALIs using the MPM. In order to do that, we focus on the chaotic orbit considered in Figs. 4 and 5. For our computations we fix $d_{0}=10^{-9}$ and evolve the chaotic orbit and its nearby orbits up to $t=200$ with using as integration time step $h=0.01$. We record the final values of the $\mbox{GALI}_{2}$ in Fig. 7(a), the $\mbox{GALI}_{3}$ in Fig. 7(b) and the $\mbox{GALI}_{4}$ in Fig. 7(c), averaged over $1000$ sets of random initial orthogonal separation vectors, varying $\tau$ in the interval between $0.01$ to $100$. We find that the dependence of the computed GALIs on the renormalization time $\tau$ can be split into two main regions. The first one corresponds to $\tau\lesssim 1$ for which the final values of the computed GALIs remain practically constant at around $5\times 10^{-5}$, $2\times 10^{-10}$ and $1.2\times 10^{-18}$ for the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ respectively. On the other hand, in the region where $\tau\gtrsim 1$, we see that increasing the values of $\tau$ leads to a roughening and an overall increase in the final values of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$. This results hints toward the appearance of large numerical errors in computing the GALI values.
@@
Figure 7: 2D HH system. Dependence of the GALI2, GALI3 and GALI4 on (a-c) the renormalization time $\tau$ and (d-f) the integration time step $h$ for the chaotic orbit considered in Figs. 4 and 5. In (a-c) the integration time step is set to $h=0.01$, while in (d-f) the renormalization time is fixed to $\tau=1$. For all computations the final time of integration is $t=200$, and random initial deviation vectors of norm $d_{0}=10^{-9}$ were used. The GALI values are obtained through averaging over $1000$ sets of random initial orthogonal deviation vectors.
@@
A similar behavior is also observed when computing the dependence of the computed GALIs on the integration time step $h$. For these numerical simulations, we kept the same set up for the numerical integration as above, setting $\tau=1$ and varying the integration time step $h$. The obtained results are depicted in Figs. 7(d)-(f) for the $\mbox{GALI}_{2}$, $\mbox{GALI}_{3}$ and $\mbox{GALI}_{4}$ respectively. We see that the dependence of the final value of the GALIs on $h$ can also be split into two main regions, separated by $h\approx 10^{-1}$. For $h\lesssim 0.1$, these GALI values are constant for all $h$, while for $h\gtrsim 0.1$ we obtain different final values of the GALIs for each time step $h$. The striking result here is that in terms of the relative energy error, $h=0.1$ corresponds to $E_{r}\approx 10^{-8}$ [Fig. 8]. Thus, we conclude that numerical simulations with relative energy errors $E_{r}\gtrsim 10^{-8}$ are in general inaccurate. Such a precision could be easily achieved using high order numerical integration schemes. It is worth pointing out the slope of the $E_{r}$ against $h$ in Fig. 8 is $4$ (this slope is indicated by the dashed black line), matching the order of the $ABA864$ symplectic integrator [see also Sect. 4.1].
@@
Figure 8: 2D HH system. Dependence of the upper bound of the time evolution of the relative energy error $E_{r}$ (41) on the integration time step $h$, for the chaotic orbit of Fig. 4, when it is integrated up to $t=200$ by the $ABA864$ symplectic integrator of order $4$. The dashed black line indicates slope $4$.
@@
What about exploring the global dynamics of a dynamical system using the GALI method with the MPM? In Fig. 9, we show colormaps of the PSS $x_{1}=0$ of Fig. 2, made of $700\times 600$ grid points. The $p_{1}$ coordinate of each initial condition on the ($x_{2},p_{2}$) phase plane of Fig. 9 is set such that $p_{1}>0$ with $\mathcal{H}=0.125$. The initial condition of each generated orbit on the $(x_{2},p_{2})$ PSS is colored according its MLE value at $t=1000$ [Fig. 9(a)] and its $\mbox{GALI}_{2}$ value at $t=500$ [Fig. 9(b)]. We also note that the presented results are obtained by averaging the indices values over $50$ configurations of sets of random orthogonal initial tangent vectors in order to mitigate their dependence on the choice of the initial conditions for the tangent vectors. Both chaos indicators display the same qualitative dynamical behavior in terms of characterizing the stability of orbits in the phase space. More specifically, Figs. 9(a-b), show the same phase plane structure, that is to say the regular regions (yellow colored points) and chaotic seas (purple colored points) are practically the same in both panels, and similar to what is revealed by the PSS of Fig. 2. In particular, we see that several small regular islands in the chaotic regions are well capture by the MPM, along with the so-called ‘sticky orbits’ at the interfaces of regular islands with chaotic seas corresponding to a mixed blend of colors (e.g., yellow purple). The difference between the two numerical simulations resides on the computational time. Indeed, it took around $t=35.7$ and $t=28.2$ CPU hours respectively for the simulations of Fig. 9(a) and Fig. 9(b) to be completed.
Figure 9: 2D HH system. Regions of different dynamical behavior on the $x_{1}=0$, $p_{1}>0$ PSS of the HH system (38) with $\mathcal{H}=0.125$, identified through (a) the MLE values at $t=1000$, and (b) the $\mbox{GALI}_{2}$ values at $t=500$. (c) Regions of different values of the time $t$ needed for the $\mbox{GALI}_{3}$ to become less than $10^{-8}$ for the same set of initial conditions as in panels (a) and (b). In all computations $h=0.01$, $\tau=1$ and $d_{0}=10^{-8}$. In each panel the average GALI values over results obtain for $50$ sets of random initial orthogonal separation vectors are shown, with initial conditions colored according to the color scale at the right side of the panel.